# SofT-GRPO: Surpassing Discrete-Token LLM Reinforcement Learning via Gumbel-Reparameterized Soft-Thinking Policy Optimization

### v1

@@ Abstract
The soft-thinking paradigm for Large Language Model (LLM) reasoning can outperform the conventional discrete-token Chain-of-Thought (CoT) reasoning in some scenarios, underscoring its research and application value. However, while the discrete-token CoT reasoning pattern can be reinforced through policy optimization algorithms such as group relative policy optimization (GRPO), extending the soft-thinking pattern with Reinforcement Learning (RL) remains challenging. This difficulty stems from the complexities of injecting stochasticity into soft-thinking tokens and updating soft-thinking policies accordingly. As a result, previous attempts to combine soft-thinking with GRPO typically underperform their discrete-token GRPO counterparts. To fully unlock the potential of soft-thinking, this paper presents a novel policy optimization algorithm, SofT-GRPO, to reinforce LLMs under the soft-thinking reasoning pattern. SofT-GRPO injects the Gumbel noise into logits, employs the Gumbel-Softmax technique to avoid soft-thinking tokens outside the pre-trained embedding space, and leverages the reparameterization trick in policy gradient. We conduct experiments across base LLMs ranging from 1.5B to 7B parameters, and results demonstrate that SofT-GRPO enables soft-thinking LLMs to slightly outperform discrete-token GRPO on Pass@1 (+0.13% on average accuracy), while exhibiting a substantial uplift on Pass@32 (+2.19% on average accuracy).  [^2] Codes and weights are available on https://github.com/zz1358m/SofT-GRPO-master
@@
(a) Reasoning with Discrete-Token CoT
(b) Reasoning with the Soft-Thinking Pattern
(c) GRPO (shao2024deepseekmath) for Discrete-Token CoT
(d) Existing Work of Soft-Thinking + GRPO (butt2025soft)
@@
Figure 1: The soft-thinking pattern (b) passes the expectation of embeddings to the next LLM step (zhang2025soft), which can surpass the conventional discrete-token CoT (a) without any fine-tuning. However, employing the GRPO algorithm (c) will boost the performance of discrete-token CoT, but existing attempts (d) of applying RLVR to soft-thinking derive inferior performances. The proposed SofT-GRPO (e) provides the first valid RLVR algorithm, which can outperform the discrete-token CoT with GRPO.
@@ 1 Introduction
Reasoning with large language models (LLMs) has demonstrated impressive versatility across diverse domains (sprague2024cot). However, most existing reasoning methods rely on generating discrete tokens, which may limit their ability to represent certain abstract concepts (zhang2025soft). In pursuit of better expressing abstract ideas and even implementing a possible latent search tree (wu2025llms), zhang2025soft presents the soft-thinking reasoning pattern. The soft-thinking reasoning pattern replaces each discrete token in the chain-of-thought (CoT) with a continuous representation: a weighted sum of d-dimensional token embeddings, computed from their output probabilities, which is then input to the next LLM decoding step. With suitable sampling techniques—such as the Gumbel-Softmax technique (wu2025llms) or the Dirichlet resampling technique (zhuang2025text)—soft-thinking can outperform conventional discrete-token CoT on a wide range of tasks, without requiring any fine-tuning (wu2025llms).
Recently, a growing body of research has focused on Reinforcement Learning with Verifiable Rewards (RLVR) approaches to further enhance the effectiveness of discrete-token CoT reasoning (wang2024reinforcement; liu2025understanding; yu2025dapo). Among these approaches, Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) has emerged as a particularly compelling framework. GRPO operates by sampling groups of CoT trajectories for each query and updating the policy to favor higher-reward trajectories, yielding substantial improvements on various benchmarks. Notably, discrete-token GRPO can consistently outperform the possible advantage brought about by applying the soft-thinking pattern on original LLMs. In contrast, initial attempts to synergize soft-thinking with GRPO typically underperform in comparison to their discrete-token counterparts (butt2025soft). As shown in 1 (b), different from discrete-token reasoning processes, tokens in soft-thinking reasoning are deterministic without randomness. So, this shortfall can be attributed to challenges of 1) introducing controllable stochasticity within soft-thinking tokens for reasoning path exploration, as well as 2): effectively updating the soft-thinking policies by exploiting high-quality samples. These challenges highlight the need for specialized policy optimization algorithms that fully unlock the latent potential of soft-thinking reasoning in LLMs.
@@
To address these challenges and unlock the promise of soft-thinking, we propose a novel policy optimization algorithm, SofT-GRPO, specifically designed to reinforce the soft-thinking reasoning ability in LLMs. As shown in Figure 1(e), in the rollout process, SofT-GRPO samples groups of soft-thinking reasoning paths by injecting sampled Gumbel noises into the output probabilities and employs the Gumbel-Softmax technique to avoid invalid inputs outside the pre-trained discrete-token embedding space. For the following soft-thinking policy updates, SofT-GRPO leverages the reparameterization trick on the Gumbel distribution. It can achieve accurate gradient estimation and accurately attribute the reward improvement to the output probability of LLMs. We conduct thorough evaluations of SofT-GRPO on three representative LLM architectures—DeepSeek-R1-Distill-Qwen-1.5B, LLaMA-3.2-3B-Instruct, and DeepSeek-R1-Distill-Qwen-7B—across five numerical reasoning benchmarks. Experimental results consistently show that SofT-GRPO-enhanced soft-thinking not only slightly surpasses discrete-token GRPO on Pass@1, but also yields substantial gains on Pass@32, highlighting the practical advantages of robust policy optimization for soft-thinking reasoning.
Our contributions can be summarized as follows:
@@
  • We introduce SofT-GRPO, a novel and powerful policy optimization algorithm designed for reinforcing the soft-thinking reasoning paradigm in LLMs. It integrates the Gumbel-Softmax technique into the group rollout process, actively obtaining diverse but valid soft-thinking reasoning paths.
@@
  • We propose an innovative gradient estimation approach via Gumbel reparameterization, enabling precise attribution of improvements to the LLM’s output probability distributions in policy optimization.
  • We conduct comprehensive experiments across LLMs of 1.5B–7B parameters on five in-domain benchmarks and three out-of-domain benchmarks, demonstrating that SofT-GRPO consistently outperforms the discrete-token GRPO baselines, especially at higher sample rates (Pass@16 or Pass@32).
@@ 2.1 Discrete-Token CoT Reasoning
Discrete-token CoT reasoning processes seek to solve a $|\boldsymbol{Q}|$-token question $\boldsymbol{Q}=(q_{1},\ldots,q_{|\boldsymbol{Q}|})$ by generating $|\boldsymbol{R}|$ reasoning CoT tokens $\boldsymbol{R}=(r_{1},\ldots,r_{|\boldsymbol{R}|})$ before outputting the answer prediction $\boldsymbol{A}=(a_{1},\ldots,a_{|\boldsymbol{A}|})$ (guo2025deepseek; sprague2024cot). All these tokens are in the discrete language domain, i.e., $\boldsymbol{Q},\boldsymbol{R},\boldsymbol{A}\subset\mathcal{T}$, where $\mathcal{T}$ represents the set of language tokens. Language reasoning tokens and answer tokens are generated with the next-token prediction (NTP) policy of LLM $\pi_{\theta}$ as follows:
@@
where $\cdot$,\ $\cdot$,\ $\cdot$] denotes concatenation.
@@
$\Bigg[\sum_{g=1}^{G}\frac{1}{\left|\boldsymbol{R}_{g}\right|+\left|\boldsymbol{A}_{g}\right|}\sum_{t=1}^{\left|\boldsymbol{R}_{g}\right|+\left|\boldsymbol{A}_{g}\right|}\Big(\min\left(p_{g,t}\hat{A}_{g},\text{clip}(p_{g,t},1-\epsilon,1+\epsilon)\hat{A}_{g}\right)-\beta D_{KL}(\pi_{\theta}||\pi_{\theta_{\text{ref}}})\Bigg]$
@@
Supervised Fine-tuning (SFT) for Discrete-Token CoT LLM Reasoning. As straightforward fine-tuning methods for LLM reasoning, SFT methods (wu2025llm; zheng2025reasoning) first collect high-quality CoT labels $\boldsymbol{R}^{*},\boldsymbol{A}^{*}]$ for each question $\boldsymbol{Q}$. Then LLMs are fine-tuned for correct predictions on each of the $|\boldsymbol{R}|+|\boldsymbol{A}|$ tokens. However, SFT methods for LLM reasoning highly rely on the quality of CoT labels, which are difficult to obtain for complex datasets. Moreover, there are also concerns about the out-of-domain generalization ability of SFT (chu2025sft).
RLVR Fine-tuning for Discrete-Token Reasoning. RLVR fine-tuning methods such as GRPO (liu2024deepseek), Dr. GRPO (liu2025understanding), DAPO (yu2025dapo), and Lite PPO (liu2025part) sample several CoTs $\boldsymbol{R},\boldsymbol{A}]$ and assign a reward to each of them based on the quality of the answers $\boldsymbol{A}$. The standard discrete-token GRPO algorithm (shao2024deepseekmath) samples $G$ CoTs for each question $\boldsymbol{Q}$ and optimizes the reasoning policy towards the CoTs with higher rewards.
The loss function of GRPO is shown in Eq. (2), where $\hat{A}_{g}$ represents the advantage function for the $g$-th CoT, the reward function $f(\boldsymbol{A}_{t})=1$ if and only if the answer $\boldsymbol{A}_{t}$ is correct, $\epsilon$ is the clipping hyperparameter, $\pi_{\theta_{\text{old}}}$ is the policy before the update, and $D_{KL}$ represents the KL-divergence of policy $\pi_{\theta}$ and a reference model policy $\pi_{\theta_{\text{ref}}}$. Employing RLVR methods, the average performance of LLMs in numerical reasoning problems can be improved without the need for labels (yue2025does).
2.2 Soft-Thinking Paradigm
Conventional discrete-token reasoning is constrained to selecting a single token from the token set $\mathcal{T}$ at each of the CoT reasoning steps. This approach may hinder the model’s ability to express certain abstract concepts that cannot be easily represented by a single deterministic token (zhang2025soft). To enhance the model’s capability to represent abstract concepts, zhang2025soft presents another paradigm called soft-thinking. The proposed soft-thinking paradigm replaces discrete-token reasoning steps $\boldsymbol{R}$ with a soft-thinking reasoning path $\boldsymbol{S}=(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{|\boldsymbol{S}|})$. Each token $\boldsymbol{s}_{i}\in\mathcal{R}^{d}$ is a real-valued vector, which is calculated as the weighted sum of token embeddings with their output probabilities. The weighted sum embeddings within the vector space of token embeddings are then fed into the next LLM step as follows:
$\boldsymbol{p}_{t}\sim\pi_{\theta}(\cdot|\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})])),$ (3)
@@ 2.2 Soft-Thinking Paradigm
where $p_{i}\in[0,1]$ is the predicted probability of token $i$ in $\boldsymbol{p}_{t}$ and $\boldsymbol{e}_{i}\in\mathcal{R}^{d}$ is the LLM embeddings of token $i$. The soft-thinking paradigm may inherently implement a multi-thread latent search tree in the serial decoding process (wu2025llms).
Based on Eq. 3, wu2025llms find that most pre-trained LLMs tend to be single-thread, and wu2025llms propose to introduce sampling strategies in the soft-thinking reasoning process for randomness, employing methods such as the Gumbel-Softmax technique (jang2016categorical) based on the output probability $\boldsymbol{p}_{t}$ as follows:
@@
where $\epsilon_{i}$ is a scaler noise sampled from the Gumbel distribution Gumbel(0,1)(0,1), and $\tau_{g}$ is the temperature of Gumbel-Softmax. Besides using Gumbel-Softmax. wu2025llms also tries to use the Dirichlet resampling technique as follows:
$(x_{1},\ldots,x_{|\mathcal{T}|})\sim f(\boldsymbol{p}_{t})=\frac{1}{B(\alpha\cdot\boldsymbol{p}_{t})}\prod_{i}^{|\mathcal{T}|}x_{i}^{\alpha\cdot p_{i}-1}$ (5)
$\qquad\boldsymbol{s}_{t}=\sum_{i=1}^{|\mathcal{T}|}x_{i}\cdot\boldsymbol{e}_{i},$
where $\alpha$ is a scaling parameter. Empirically, cooperated with the Gumbel-Softmax technique for randomness, the soft-thinking pattern can outperform conventional discrete-token CoT on a broad range of tasks, including numerical reasoning, code reasoning, and scientific reasoning, without requiring any fine-tuning. (wu2025llms).
Figure 2: The pipeline of the proposed SofT-GRPO algorithm. In training with a Query $\boldsymbol{Q}$, the SofT-GRPO first generates a group of $G$ soft-thinking reasoning paths with Gumbel noises and the Gumbel-Softmax technique (jang2016categorical). We transmit the value $g^{\prime}_{i}$ and $y^{\prime}_{i}$ for the loss calculation afterward. Then, we reconstruct the soft-thinking input. Finally, we update the soft-thinking policy with the off-policy REINFORCE (williams1992simple) algorithm, optimizing the soft-thinking reasoning tokens with Gumbel reparameterization.
2.3 Attempts of RLVR on Soft-Thinking
However, after being boosted with the RLVR fine-tuning, discrete-token CoT will turn to clearly outperform the soft-thinking reasoning. So, butt2025soft tries to similarly improve the performance of soft-thinking using GRPO. butt2025soft adds Gaussian noise on the input $s_{t}$ in Eq. (3) as follows:

### v2

@@ Abstract
The soft-thinking reasoning paradigm has demonstrated superior performance over traditional discrete-token Chain-of-Thought (CoT) reasoning in various scenarios. However, while discrete-token CoT reasoning can be reinforced through advanced reinforcement learning with verifiable rewards (RLVR) techniques such as group relative policy optimization (GRPO), extending the soft-thinking reasoning with such strong techniques remains challenging. This difficulty stems from the complexities of injecting stochasticity into soft-thinking tokens and updating soft-thinking policies accordingly. As a result, previous attempts to combine soft-thinking with RLVR typically underperform their discrete-token RLVR counterparts. To fully unlock the potential of soft-thinking, this paper presents a powerful policy optimization algorithm, SofT-GRPO. It injects the Gumbel noise into token probabilities with Gumbel-Softmax for controllable stochasticity, and leverages the Gumbel reparameterization trick to achieve accurate credit assignment to LLM soft-thinking policies. We conduct experiments over LLMs ranging from 1.5B to 7B parameters, where SofT-GRPO enables LLMs with soft-thinking to slightly outperform discrete-token CoT GRPO on Pass@1 (+0.13% on average accuracy), and brings a substantial uplift on Pass@32 (+2.19% on average) [^2] The codes are available at https://github.com/zz1358m/SofT-GRPO-master.
@@
(a) LLM Reasoning with Discrete-Token CoT
(b) LLM Reasoning with Soft-Thinking (Zhang et al., 2025b)
(c) GRPO (Shao et al., 2024) for Discrete-Token CoT
(d) Existing Work of Soft-Thinking + GRPO (Butt et al., 2025)
@@
Figure 1: The soft-thinking pattern (b) passes the expectation of token embeddings to the next LLM step (Zhang et al., 2025b), which can surpass the conventional discrete-token CoT (a) without any fine-tuning. However, employing the GRPO algorithm (c) will boost the performance of discrete-token CoT over soft-thinking, and existing attempts (d) of applying RLVR to soft-thinking derive inferior performances. The proposed SofT-GRPO (e) provides the first powerful RLVR algorithm, which can outperform the discrete-token CoT with GRPO on Pass@1, especially Pass@K.
@@ 1 Introduction
Reasoning with large language models (LLMs) has demonstrated impressive versatility in diverse domains (Sprague et al., 2024). However, most existing reasoning methods rely on the generation of discrete language tokens, which will limit their ability to represent certain abstract concepts that cannot be accurately represented by a discrete token (Zhang et al., 2025b). In pursuit of better expression of abstract ideas, Zhang et al. (2025b) presents the soft-thinking reasoning pattern. As shown in Figure 1 (b), it replaces each discrete token in the chain-of-thought (CoT) with a continuous representation: the weighted sum of $d$-dimensional token embeddings, taking the output probabilities as the weights. With suitable sampling techniques—such as with the Gumbel-Softmax technique (Wu et al., 2025a) or the Dirichlet resampling technique (Zhuang et al., 2025)—soft-thinking can outperform discrete-token CoT on a wide range of tasks without requiring any fine-tuning (Wu et al., 2025a).
Recently, reinforcement learning with verifiable rewards (RLVR) approaches have become standard for improving discrete-token CoT reasoning (Wang et al., 2024; Liu et al., 2025b; Yu et al., 2025). The most representative RLVR method, Group Relative Policy Optimization (GRPO; Figure 1 (c)) (Shao et al., 2024), samples groups of CoT trajectories per query and reinforces the policy toward higher-reward examples. Notably, RLVR fine-tuning can yield discrete-token CoT performance that surpasses that of soft-thinking reasoning from the original, unfine-tuned LLMs. Moreover, after discrete-token GRPO, the model’s soft-thinking ability often does not improve as substantially as its discrete-token reasoning ability.
However, extending RLVR methods to soft-thinking reasoning remains challenging (Jain and Rappazzo, 2025). As illustrated in Figure 1 (b), in contrast to the stochastic nature of sampling-based discrete-token CoTs, soft-thinking reasoning produces deterministic soft-tokens, lacking both randomness and gradients for RLVR. So, extending RLVR to soft-thinking should (1) introduce controllable stochasticity to adequately explore diverse soft-thinking paths, and (2) enable unbiased policy updates that properly assign the credits of high-reward soft-thinking samples to LLM logits. Previous work (Butt et al., 2025) (see Figure 1 (d)) attempts to address stochasticity by injecting Gaussian noise into soft tokens and applying a Gaussian reparameterization trick for RLVR gradients, but such approaches have struggled to match the performance of discrete-token RLVR, largely due to difficulties in attributing RLVR advantages to logits.
@@
To address these challenges and fully unlock the LLM reasoning ability by soft-thinking, we propose a powerful policy optimization algorithm, SofT-GRPO. As shown in Figure 1(e), to introduce controllable stochasticity in the rollout process, SofT-GRPO samples groups of soft-thinking reasoning paths by injecting Gumbel noises into the output token probabilities and employs the Gumbel-Softmax technique to avoid invalid soft-tokens outside the pre-trained discrete-token embedding space. To accurately assign the reward credits to the output probability of LLMs in soft-thinking policy updates, SofT-GRPO leverages the reparameterization trick on the Gumbel distribution. We perform thorough evaluations of SofT-GRPO on three representative LLMs from 1.5B to 7B parameters and across five numerical reasoning benchmarks. Experimental results show that SofT-GRPO-enhanced soft-thinking not only surpasses discrete-token GRPO on the averaged Pass@1 but also yields substantial gains on Pass@32, highlighting the practical advantages of robust policy optimization for soft-thinking reasoning. Our contributions are as follows:
@@
  • We present SofT-GRPO, a powerful RLVR algorithm designed with the soft-thinking reasoning paradigm. It adopts the Gumbel-Softmax technique into the group rollout process, actively obtaining diverse but valid soft-thinking reasoning paths.
@@
  • SofT-GRPO proposes a novel gradient estimation approach via Gumbel reparameterization, first achieving precise attribution of improvements to the LLM’s output probability distributions in policy optimization.
  • Over eight in-domain and out-of-domain benchmarks, SofT-GRPO improves the LLM’s reasoning ability with soft-thinking, outperforming the discrete-token GRPO, especially at higher sample rates (Pass@16 or Pass@32).
@@ 2.1 Discrete-Token CoT Reasoning
To solve a $|\boldsymbol{Q}|$-token question $\boldsymbol{Q}=(q_{1},\ldots,q_{|\boldsymbol{Q}|})$, the discrete-token CoT reasoning process generates $|\boldsymbol{R}|$ reasoning CoT tokens $\boldsymbol{R}=(r_{1},\ldots,r_{|\boldsymbol{R}|})$ before predicting the answer $\boldsymbol{A}=(a_{1},\ldots,a_{|\boldsymbol{A}|})$ (Guo et al., 2025; Sprague et al., 2024). Each of $\boldsymbol{Q}$, $\boldsymbol{R}$, and $\boldsymbol{A}$ is a token sequence over the vocabulary $\mathcal{T}$ (i.e., $\boldsymbol{Q},\boldsymbol{R},\boldsymbol{A}\in\mathcal{T}^{*}$), where $\mathcal{T}$ is the set of all possible language tokens. Language reasoning tokens and answer tokens are generated with the next-token prediction (NTP) policy of LLM $\pi_{\theta}$ as follows:
@@
where $\cdot$,\ $\cdot$] and $\cdot$,\ $\cdot$,\ $\cdot$] denotes concatenation.
@@
$\Bigg[\sum_{g=1}^{G}\frac{1}{\left|\boldsymbol{R}_{g}\right|+\left|\boldsymbol{A}_{g}\right|}\sum_{t=1}^{\left|\boldsymbol{R}_{g}\right|+\left|\boldsymbol{A}_{g}\right|}\Big(\min\left(p_{g,t}\hat{A}_{g},\text{clip}(p_{g,t},1-\epsilon,1+\epsilon)\hat{A}_{g}\right)-\beta\mathbb{D}_{KL}(\pi_{\theta}||\pi_{\theta_{\text{ref}}})\Bigg]$
@@
Supervised Fine-tuning (SFT) for Discrete-Token Reasoning. As straightforward fine-tuning methods for LLM reasoning, the SFT methods (Wu et al., 2025b; Zheng and Lee, 2025) collect high-quality CoT labels for each question and then fine-tune LLMs for correct predictions. However, SFT methods highly rely on the quality of CoT labels, which are hard to obtain for complex datasets. Moreover, there are also concerns about the out-of-domain generalization ability of SFT (Chu et al., 2025).
RLVR Fine-tuning for Discrete-Token Reasoning. RLVR fine-tuning methods such as GRPO (Liu et al., 2024), Dr. GRPO (Liu et al., 2025b), and DAPO (Yu et al., 2025) sample several CoTs $\boldsymbol{R},\boldsymbol{A}]$ and assign a reward to each of them based on the quality of the answers $\boldsymbol{A}$. The standard discrete-token GRPO algorithm (Shao et al., 2024) samples $G$ CoTs for each question $\boldsymbol{Q}$ with sampling policy $\pi_{\theta_{\text{old}}}$ and optimizes the reasoning policy towards CoTs with higher rewards. Its loss function is shown in Eq. (2), where $\hat{A}_{g}$ represents the advantage function for the $g$-th CoT, the reward function $f(\boldsymbol{A}_{g})=1$ if and only if the answer $\boldsymbol{A}_{g}$ is correct, $\epsilon$ is the clipping hyperparameter, $p_{g,t}$ is calculated in an off-policy way for both the current policy $\pi_{\theta}$ and the sampling policy $\pi_{\theta_{\text{old}}}$, and $\mathbb{D}_{KL}$ represents the KL-divergence of $\pi_{\theta}$ and a reference policy $\pi_{\theta_{\text{ref}}}$. RLVR methods can improve the discrete-token reasoning performance of LLMs across a wide range of reasoning problems without the need for CoT labels. However, for each CoT in each LLM step, discrete-token GRPO updates the probability of only one token based on the reward. As mentioned in Yue et al. (2025), making RLVR does not really incentivize reasoning capacity in LLMs beyond the base model.
2.2 Soft-Thinking Reasoning Paradigm
At each of the CoT reasoning steps, conventional discrete-token reasoning is constrained to selecting one token from the token set $\mathcal{T}$, which may hinder the model’s ability to express certain abstract concepts that cannot be easily represented by a single deterministic token (Zhang et al., 2025b), thus undermining the LLM’s expression ability in reasoning. To represent abstract concepts, Zhang et al. (2025b) presents the soft-thinking paradigm, which replaces discrete-token reasoning steps $\boldsymbol{R}$ with a soft-thinking reasoning path $\boldsymbol{S}=(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{|\boldsymbol{S}|})$. Each token $\boldsymbol{s}_{i}\in\mathcal{R}^{d}$ (noted soft token) is a real vector, which is the sum of token embeddings weighted by their respective output probabilities. They are then fed into the next LLM step as follows:
$\boldsymbol{p}_{t}=\pi_{\theta}(\cdot|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})])),$ (3)
@@ 2.2 Soft-Thinking Reasoning Paradigm
where $p_{i}\in[0,1]$ is the predicted probability of token $i$ in $\boldsymbol{p}_{t}$ and $\boldsymbol{e}_{i}\in\mathcal{R}^{d}$ is the LLM embeddings of token $i$. As detailed in Appendix A.3, the soft tokens lie in the convex hull of token embeddings, making it possible to maintain validity while transmitting more information.
Based on Eq. (3), Wu et al. (2025a) find that most pre-trained LLMs tend to be single-threaded and the vanilla soft-thinking reasoning may lead to greedy feedback loop that suppresses alternative reasoning paths and undermines the benefits of transmitting more informative soft-tokens. So, they propose to introduce sampling strategies in the soft-thinking reasoning process for randomness, employing Gumbel noises and Gumbel-Softmax technique (Maddison et al., 2016; Jang et al., 2016) based on the output probability $\boldsymbol{p}_{t}$ as follows:
@@
where $\epsilon_{i}$ is a scalar noise sampled from the Gumbel distribution Gumbel(0,1)(0,1), and $\tau_{g}$ is the temperature of Gumbel-Softmax. Besides Gumbel-Softmax, Wu et al. (2025a) also tries to use the Dirichlet resampling technique as follows:
$(x_{1},\ldots,x_{|\mathcal{T}|})\sim\text{Dirichlet}(\alpha,\boldsymbol{p}_{t}),\quad\boldsymbol{s}_{t}=\sum_{i=1}^{|\mathcal{T}|}x_{i}\cdot\boldsymbol{e}_{i},$ (5)
where $\alpha$ is a scaling parameter. Empirically, incorporated with Gumbel noises and the Gumbel-Softmax technique, the soft-thinking pattern can outperform conventional discrete-token CoT on a broad range of tasks, including numerical, code, and scientific reasoning, without requiring any fine-tuning. (Wu et al., 2025a).
Figure 2: The pipeline of the proposed SofT-GRPO algorithm. In training with each Query $\boldsymbol{Q}$, the SofT-GRPO first generates a group of $G$ soft-thinking reasoning paths with Gumbel noises and the Gumbel-Softmax technique (Wu et al., 2025a). We transmit the value $g^{\prime}_{i}$ and $y^{\prime}_{i}$ for the loss calculation afterward. Then, we reconstruct the soft-thinking input. Finally, we update the soft-thinking policy with the off-policy proximal policy optimization (Schulman et al., 2017), optimizing soft-thinking policies with Gumbel reparameterization.
2.3 Related Work on RLVR Soft-Thinking
However, after being boosted with the RLVR fine-tuning, discrete-token reasoning will clearly outperform soft-thinking. So, Butt et al. (2025) tries to similarly improve the soft-thinking reasoning ability using GRPO by adding Gaussian noise on the soft-token $\boldsymbol{s}_{t}$ in Eq. (3) as follows: